this is a two part question. A and B please it is one question only!!5) Use the derivative of y = f(x) = x3 + 4x2 – x - 8 to a) determine the slope of the tangent line on the graph of y = f(x) ab) give the equation of the tangent line to the graph of y = f(x) at x = -3.THANK YOU
Answer #1

as f(x) = 23 5) = ré + 43-2-8 f(x) = 3x²+8x-1 (@ the slope of the tangent on the graph y=f(x) at the point a=-3 is f(-3) N